Grade 8Magical Mathematics[Interesting Approach]

​the number of ways in which 85 can be written as a sum of two squares

dheeraj
9 Years agoGrade 8

1 Answer

Ritika Das
ApprovedApproved Tutor Answer9 Years ago
The number of ways in which 85 can be written as a sum of two squares is by two ways. They are as follows: – 22+92=4+81=85
62+72=36+49=85
Please approve my answer.
Regards.